# Prompt-prompted Mixture of Experts for Efficient LLM Generation

### v1

@@ Abstract
With the development of transformer-based large language models (LLMs), they have been applied to many fields due to their remarkable utility, but this comes at a considerable computational cost at deployment. Fortunately, some methods such as pruning or constructing a mixture of experts (MoE) aim at exploiting sparsity in transformer feedforward (FF) blocks to gain boosts in speed and reduction in memory requirements. However, these techniques can be very costly and inflexible in practice, as they often require training or are restricted to specific types of architectures. To address this, we introduce GRIFFIN, a novel training-free MoE that selects unique FF experts at the sequence level for efficient generation across a plethora of LLMs with different non-ReLU activation functions. This is possible due to a critical observation that many trained LLMs naturally produce highly structured FF activation patterns within a sequence, which we call flocking. Despite our method’s simplicity, we show with 50% of the FF parameters, GRIFFIN maintains the original model’s performance with little to no degradation on a variety of classification and generation tasks, all while improving latency (e.g. 1.25$\times$ speed-up in Llama 2 13B on an NVIDIA L40). Code will be available at https://github.com/hdong920/GRIFFIN.
@@ 1 Introduction
Transformers [VSP+17] have demonstrated incredible capabilities across a plethora of domains [LWLQ22, KNH+22, NBZ+23]. Their large language model (LLM) successors [TMS+23, TAB+23, JSM+23, JSR+24, TMH+24, Ant24] have pushed the bar higher, but these behemoths have become performative at the price of enormous amounts of computation and memory demands. One significant contributor is the model size itself. Not only is storage an issue, model layers tend to be wide and plenty, slowing down inference. Moreover, given the existence of sparse structures in LLMs, especially in feedforward (FF) blocks [GSBL20, DLBZ22, LYB+22, LWD+23], these models waste computation on intermediate features with little to no impact on the final result. For instance, it has been observed that in OPT-175B [ZRG+22], fewer than 5% of neurons in FF blocks have nonzero values per token [LWD+23], meaning 95% of the compute in each FF block is wasted. Usually consisting of around two-thirds of the parameters in an LLM, FF blocks can be serious memory and compute bottlenecks. These inefficiencies are highly problematic in latency-sensitive scenarios like in chatbots and autonomous vehicles.
There have been many methods to exploit sparsity in LLMs for efficiency gains, such as pruning model weights and constructing mixtures of experts (MoEs). Pruning removes low-impact pre-trained weights to reduce storage, yet this often does not translate to real speed-ups in practice, unless the pruning is done in a hardware-friendly manner which typically causes greater performance deterioration. MoEs better preserve the original performance by adaptively selecting subsets of the model to use per input but also come with drawbacks. Unless the model has been trained in this fashion [FZS22, JSR+24], it will need to learn a cheap yet effective gating function (expert selection mechanism) and sometimes even require full fine tuning. Perhaps an even bigger weakness of many of these methods is the inability to effectively carry over to pre-trained LLMs with non-ReLU activations. We seek to overcome these challenges with MoEs:
@@
Figure 1: Relative FF activation magnitudes of the first 512 features and tokens across a sequence from PG-19 [RPJ+19, GBB+20] in layer 10 of Llama 2 7B (left) and Gemma 7B (right). These heatmaps show flocking, where relative activation magnitudes are shared within a sequence. More examples in Appendix B.
@@
    Model & Activation Function Diversity: Thorough experimentation demonstrates the efficacy of GRIFFIN on numerous models, including Llama 2 [TMS+23], Gemma [TMH+24], Mistral [JSM+23], OPT [ZRG+22], and ReluLlama [Tea23]. Together, the tested activation functions include ReLU, SwiGLU, GEGLU, and ReGLU [Sha20].
@@ 2 Background
The observation that transformer FF blocks produce sparse activations is not new [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23]. In ReLU-based LLMs like OPT [ZRG+22], the activations can be exceptionally sparse and become more apparent for larger models [LWD+23]. As more models use non-sparse activation functions like GLU variants [Sha20], it is difficult for neurons to have no contribution to the output since these functions do not have an interval that maps to zero. Without exact sparsity, the efficacy of these methods becomes limited. As such, this has ushered a wave of models that are either adapted from available models [ZLL+21, LLLC21, MAM+23, ZBC+24, JSR+24] or trained from scratch [FZS22] which can produce activations that are exactly zero with little to no performance loss. Even so, these methods require considerable amounts of computational resources.
@@
Pruning [LDS89] is another sparsity-guided way to tackle compute and memory bottlenecks of models. Previously, the common method would be some variation of iteratively rounding weights down to zero based on some score and retraining to recover any lost performance [FC18, BGOFG20, LGW+21, LZH+24]. While this can result in most parameters being pruned, this method comes with a few issues. First, with the increasing scale of LLMs, retraining becomes impractical for most. Fortunately, cheap methods to effectively prune LLMs have been developed [FA23, SLBK23, JLC+23, DKK+24]. The second issue is that unless pruning is done in a structured manner [XZC22, SWSL23, MFW23, LYZ+23, XGZC23], it is difficult to see real computational savings, yet structured pruning often leads to much more severe performance degradation. Third, pruning enforces sparsity to be static which can be strong assumption since FF blocks are widely believed to contain the model’s memory [GSBL20].
@@
Making sparsity more dynamic has motivated the design of mixture-of-experts (MoEs) [JJNH91] to avoid computing low-impact features in trained models at varying granularities [ZLL+21, LWD+23, PSBW23, CIS23, AMB+23, YYB+24, ZBC+24]. The main idea of MoEs is to use a gating function to identify a small subset of neurons that will be used to compute the layer’s output, an adaptive form of pruning. In the ideal case, all active neurons are selected and inactive neurons are ignored for each input. However, current methods either require training or rely on ReLU activation functions. Our method has the best of both worlds: it is training-free and effective on a variety of activation functions.
@@ 5.1 Performance
Using various models, we evaluate on several generation and classification tasks. For generation, we evaluate on XSum [NCL18], CNN/DailyMail [NZG+16], COQA [RCM19], and SCROLLS QASPER [DLB+21, SSI+22]. For classification, we evaluate on HellaSwag [ZHB+19], PIQA [BZB+20], COPA [RBG11], ARC-Easy/Challenge [CCE+18], and BoolQ [CLC+19]. With the exception of XSum and CNN/DailyMail, we use LM Evaluation Harness for our experiments [GTA+23]. Aside from comparing with the original LLM, we also compare GRIFFIN with a static sequence-level MoE based on neuron magnitudes. Similar to neuron magnitude pruning, this baseline selects experts based on neuron magnitudes in $\bm{W}_{1}$ for the generation phase but uses the entire FF blocks for prompting like GRIFFIN. In the case of GLU variants, the neuron-wise norms of $\bm{W}_{1}$ and $\bm{W}_{g}$ are elementwise multiplied to produce the pruning metric. As we will see, this straightforward baseline achieves great classification results but falters for generation.
@@
We start with a look into the relationship between the sparsity levels and performance degradation. This translates to varying $k$ when we select the top-$k$ of our statistic $\bm{s}$. To compare the performance degradation across multiple tasks, we plot the ratio of the final performance metrics between GRIFFIN and the full model in Figure 4. We see most of the performance is preserved at 50% FF sparsity in Llama 2 7B, Gemma 7B, and Mistral 7B. Different tasks have different tipping points where performance sharply drops, which may be related to the difficulty of the task [YJL+24].

### v2

@@ Abstract
With the development of transformer-based large language models (LLMs), they have been applied to many fields due to their remarkable utility, but this comes at a considerable computational cost at deployment. Fortunately, some methods such as pruning or constructing a mixture of experts (MoE) aim at exploiting sparsity in transformer feedforward (FF) blocks to gain boosts in speed and reduction in memory requirements. However, these techniques can be very costly and inflexible in practice, as they often require training or are restricted to specific types of architectures. To address this, we introduce GRIFFIN, a novel training-free MoE that selects unique FF experts at the sequence level for efficient generation across a plethora of LLMs with different non-ReLU activation functions. This is possible due to a critical observation that many trained LLMs naturally produce highly structured FF activation patterns within a sequence, which we call flocking. Despite our method’s simplicity, we show with 50% of the FF parameters, GRIFFIN maintains the original model’s performance with little to no degradation on a variety of classification and generation tasks, all while improving latency (e.g. 1.25$\times$ speed-up in Llama 2 13B on an NVIDIA L40). Code is available at https://github.com/hdong920/GRIFFIN.
@@ 1 Introduction
Transformers [${}^{+}$17] have demonstrated incredible capabilities across a plethora of domains [LWLQ22, ${}^{+}$22, ${}^{+}$23]. Their large language model (LLM) successors [${}^{+}$23, ${}^{+}$23, ${}^{+}$23, ${}^{+}$24, ${}^{+}$24, Ant24] have pushed the bar higher, but these behemoths have become performative at the price of enormous amounts of computation and memory demands. One significant contributor is the model size itself. Not only is storage an issue, model layers tend to be wide and plenty, slowing down inference. Moreover, given the existence of sparse structures in LLMs, especially in feedforward (FF) blocks [GSBL20, DLBZ22, ${}^{+}$22, ${}^{+}$23], these models waste computation on intermediate features with little to no impact on the final result. For instance, it has been observed that in OPT-175B [${}^{+}$22], fewer than 5% of neurons in FF blocks have nonzero values per token [${}^{+}$23], meaning 95% of the compute in each FF block is wasted. Usually consisting of around two-thirds of the parameters in an LLM, FF blocks can be serious memory and compute bottlenecks. These inefficiencies are highly problematic in latency-sensitive scenarios like in chatbots and autonomous vehicles.
There have been many methods to exploit sparsity in LLMs for efficiency gains, such as pruning model weights and constructing mixtures of experts (MoEs). Pruning removes low-impact pre-trained weights to reduce storage, yet this often does not translate to real speed-ups in practice, unless the pruning is done in a hardware-friendly manner which typically causes greater performance deterioration. MoEs better preserve the original performance by adaptively selecting subsets of the model to use per input but also come with drawbacks. Unless the model has been trained in this fashion [FZS22, ${}^{+}$24], it will need to learn a cheap yet effective gating function (expert selection mechanism) and sometimes even require full fine tuning. Perhaps an even bigger weakness of many of these methods is the inability to effectively carry over to pre-trained LLMs with non-ReLU activations. We seek to overcome these challenges with MoEs:
@@
Figure 1: Relative FF activation magnitudes of the first 512 features and tokens across a sequence from PG-19 [${}^{+}$19, ${}^{+}$20] in layer 10 of Llama 2 7B (left) and Gemma 7B (right). These heatmaps show flocking, where relative activation magnitudes are shared within a sequence. More examples in Appendix B.
@@
    Model & Activation Function Diversity: Thorough experimentation demonstrates the efficacy of GRIFFIN on numerous models, including Llama 2 [${}^{+}$23], Gemma [${}^{+}$24], Mistral [${}^{+}$23], OPT [${}^{+}$22], and ReluLlama [Tea23]. Together, the tested activation functions include ReLU, SwiGLU, GEGLU, and ReGLU [Sha20].
@@ 2 Background
The observation that transformer FF blocks produce sparse activations is not new [GSBL20, DLBZ22, ${}^{+}$22, DCC23, ${}^{+}$23]. In ReLU-based LLMs like OPT [${}^{+}$22], the activations can be exceptionally sparse and become more apparent for larger models [${}^{+}$23]. As more models use non-sparse activation functions like GLU variants [Sha20], it is difficult for neurons to have no contribution to the output since these functions do not have an interval that maps to zero. Without exact sparsity, the efficacy of these methods becomes limited. As such, this has ushered a wave of models that are either adapted from available models [${}^{+}$21, LLLC21, ${}^{+}$23, ${}^{+}$24, ${}^{+}$24] or trained from scratch [FZS22] which can produce activations that are exactly zero with little to no performance loss. Even so, these methods require considerable amounts of computational resources.
@@
Pruning [LDS89] is another sparsity-guided way to tackle compute and memory bottlenecks of models. Previously, the common method would be some variation of iteratively rounding weights down to zero based on some score and retraining to recover any lost performance [FC18, BGOFG20, ${}^{+}$21, ${}^{+}$24]. While this can result in most parameters being pruned, this method comes with a few issues. First, with the increasing scale of LLMs, retraining becomes impractical for most. Fortunately, cheap methods to effectively prune LLMs have been developed [FA23, SLBK23, ${}^{+}$23, ${}^{+}$24]. The second issue is that unless pruning is done in a structured manner [XZC22, SWSL23, MFW23, ${}^{+}$23, XGZC23], it is difficult to see real computational savings, yet structured pruning often leads to much more severe performance degradation. Third, pruning enforces sparsity to be static which can be strong assumption since FF blocks are widely believed to contain the model’s memory [GSBL20].
@@
Making sparsity more dynamic has motivated the design of mixture-of-experts (MoEs) [JJNH91] to avoid computing low-impact features in trained models at varying granularities [${}^{+}$21, ${}^{+}$23, PSBW23, CIS23, ${}^{+}$23, ${}^{+}$24, ${}^{+}$24]. The main idea of MoEs is to use a gating function to identify a small subset of neurons that will be used to compute the layer’s output, an adaptive form of pruning. In the ideal case, all active neurons are selected and inactive neurons are ignored for each input. However, current methods either require training or rely on ReLU activation functions. Our method has the best of both worlds: it is training-free and effective on a variety of activation functions.
@@ 5.1 Performance
Using various models, we evaluate on several generation and classification tasks. For generation, we evaluate on XSum [NCL18], CNN/DailyMail [${}^{+}$16], COQA [RCM19], and SCROLLS QASPER [${}^{+}$21, ${}^{+}$22]. For classification, we evaluate on HellaSwag [${}^{+}$19], PIQA [${}^{+}$20], COPA [RBG11], ARC-Easy/Challenge [${}^{+}$18], and BoolQ [${}^{+}$19]. With the exception of XSum and CNN/DailyMail, we use LM Evaluation Harness for our experiments [${}^{+}$23]. Aside from comparing with the original LLM, we also compare GRIFFIN with a static sequence-level MoE based on neuron magnitudes. Similar to neuron magnitude pruning, this baseline selects experts based on neuron magnitudes in $\bm{W}_{1}$ for the generation phase but uses the entire FF blocks for prompting like GRIFFIN. In the case of GLU variants, the neuron-wise norms of $\bm{W}_{1}$ and $\bm{W}_{g}$ are elementwise multiplied to produce the pruning metric. As we will see, this straightforward baseline achieves great classification results but falters for generation.
@@
We start with a look into the relationship between the sparsity levels and performance degradation. This translates to varying $k$ when we select the top-$k$ of our statistic $\bm{s}$. To compare the performance degradation across multiple tasks, we plot the ratio of the final performance metrics between GRIFFIN and the full model in Figure 4. We see most of the performance is preserved at 50% FF sparsity in Llama 2 7B, Gemma 7B, and Mistral 7B. Different tasks have different tipping points where performance sharply drops, which may be related to the difficulty of the task [${}^{+}$24].